# Repetition Neurons: How Do Language Models Produce Repetitions?

### v1

@@ 1 Introduction
Since the advent of LLMs such as GPT-3 Brown et al. (2020), text generation has garnered significant attention from a broad range of users. Although text generation has been actively studied, the issue of repetition remains one of the fundamental challenges Li et al. (2023a); Ivgi et al. (2024). Specifically, repetition becomes critically problematic in a greedy generation, often used when reproducibility must be guaranteed Song et al. (2024).
Many researchers have attempted to address this problem from the perspectives of analyzing repetition phenomena Fu et al. (2021); Xu et al. (2022) and developing techniques to mitigate the repetition issue Keskar et al. (2019); Shirai et al. (2021); Zhu et al. (2023); Li et al. (2023a). However, the internal mechanisms of generative models when they produce repetitive outputs remain insufficiently explored Vaidya et al. (2023).
This paper focuses on the neurons of Transformer language models Vaswani et al. (2017); Geva et al. (2021); Dai et al. (2022); Chen et al. (2024) that detect repetition in inputs and trigger repetitive outputs in text generation. We refer to these neurons as “repetition neurons”. We hypothesize that repetition neurons are a type of “skill neuron” Radford et al. (2017); Wang et al. (2022) that prompts the model to generate repetition as a task of copying the previous context  [^1] We discussed the relationship between our work and most recent research Wang et al. (2024) in Appendix A., akin to “task vectors” Hendel et al. (2023) found in in-context learning Brown et al. (2020); Yan et al. (2024).
@@
We inspected the repetition neurons in three English and one Japanese pre-trained language model. Experimental results show that repetition neurons are located in both intermediate and final layers (§3.2). We further demonstrate that deactivating repetition neurons suppress the output probabilities of repeated tokens (§4.1) while activating them increases these probabilities (§4.2) [^2] We publish the codes for experiments: https://github.com/tatHi/repetition_neuron.
@@ 3.2 Observation of Repetition Neuron
Figure 3 presents the distribution of repetition neurons across different layers. The last layer contains the largest number of repetition neurons in all models, while a secondary peak appears in the intermediate layers. This suggests the existence of two types of repetition neurons: those that detect repeating patterns in the intermediate layers and those that drive the model to replicate previous contexts in the uppermost layer. The presence of repetition neurons in both the final and intermediate layers aligns with previous findings that task-specific neurons tend to reside in higher layers Wang et al. (2022), and task-related parameters and hidden states are often found in intermediate layers Hendel et al. (2023); Merullo et al. (2024). Figure 7(a) and 7(b) in Appendix E show that the location patterns of the repetition neurons remain consistent across variations in hyperparameters $|X|$ and $r$.
@@ 4 Intervention to Repetition Neurons
The latest trend in design for the kitchen sink drain is the use of a stainless steel sink drain. This is a great way to add a touch of class to your kitchen. Stainless steel sinks are also very durable and easy to clean. The stainless steel sink drain is a great way to add a touch of class to your kitchen. It is also very durable and easy to clean. What is a stainless steel sink drain? A stainless steel sink drain is a type of sink drain that is made from stainless steel. Stainless steel is a type of metal that is resistant to corrosion and rust. The latest trend in design for the kitchen sink drain is the use of a stainless steel sink drain. This is a great way to add a touch of class to your kitchen. Stainless steel sinks are also very durable and easy to clean. The stainless steel sink drain is a great way to add a touch of class to your kitchen. Stainless steel sinks are also very durable and easy to clean. The stainless steel sink drain is a great way to add a touch of class to your kitchen. Stainless steel sinks are also very durable and easy to clean. The
Table 1: The example of generation by Gemma-2B with and without intervention to the repetition neurons. The range after the intervention starts is highlighted in bold and the detected repetition is highlighted with underline.
@@ 4.1 Preventing Repetition
Result: Figure 4 shows the number of samples containing repetition after deactivating varying numbers of repetition neurons (solid lines) compared to randomly selected neurons (dashed lines). As the figure demonstrates, deactivating the repetition neurons effectively reduces the number of samples with repetition compared to deactivating randomly selected neurons. This result confirms that the repetition neurons identified by our method are indeed responsible for causing the repetition problem. We observed that deactivating repetition neurons reduces the number of samples with repetition by up to 25% (and by as much as 35% with optimal hyperparameter settings, as shown in Figures 8(a) and 8(b)). This suggests that roughly 30% of the repetition problem can be attributed to the repetition neurons. Table 3 in §C provides an example where repetition was successfully suppressed, illustrating that the generated text remains grammatically coherent despite neuron intervention. Besides, the perplexity is not largely damaged by deactivating the repetition neurons, as shown in Figure 6(a), which supports the coherency of the performance quantitatively. This confirms that the repetition neurons are specifically responsible for triggering repetition.
@@ 5 Conclusion
  • Beyond the neuron-based analysis, can we find the specific circuit in the LLMs’ calculation when outputting the repetitive texts?
@@ Appendix D Perplexity with Intervention
Figure 6(a) and 6(b) show the changes in perplexity based on different numbers of intervened repetition neurons. We used the test split of WikiText-2 for all models, including LLM-jp-1.8B, which was fine-tuned on Japanese corpora. As shown in Figure 6(a), the performance degradation caused by deactivating repetition neurons is relatively moderate. Although the impact on perplexity is smaller when intervening on randomly sampled neurons, considering that GPT-2’s perplexity on the same corpus was 18.34 Radford et al. (2019), the degradation caused by deactivating repetition neurons is acceptable. This suggests that repetition neurons do not significantly affect the generation of normal texts that do not contain repetition.
@@ Appendix E Ablation Study
Figure 7(a) and 7(b) show the location of the repetition neuron on the Transformer layers (§3.2). As shown in the figure, the size of the dataset $|X|$ does not have large effect on the distribution, which means we can obtain the similar set of repetition neurons both with smaller and larger sizes of datasets. On the other hand, $r$ has an effect on the distribution to some degree. For the case of Gemma-2B, we can obtain roughly a similar tendency with $15\leq r$.
@@
Figure 8(a) and 8(b) show the difference in the performance for the experiment about deactivating repetition neurons (§4.1). The experimental result indicates that we can obtain the more reduction effect with the larger number of deactivated repetition neurons. In contrast, the larger $r$ cause the decrease of the effect to reduce the repetition with larger number of deactivated neurons. Figure 8(b) suggests that $r=10$ or $r=15$ leads to the largest effect on controlling the repetitive generation.
@@
The experimental results for the effect of various hyperparameters on the setting with activating repetition neurons in Figure 9(a) and 9(b) show similar trends. The larger $|X|$ leads to the larger number of repetitive texts while $r=10$ or $r=15$ has the largest effect on the repetitive generation. These results could be an important clue to investigate the relation between the length of repetitive phrases and the repetition neurons in the future work.
@@ Appendix F Comparison between Two Model Sizes
Figure 11 and 12 show the experimental results of the two experiments with deactivating and activating the repetition neurons, respectively. The results of Figure 11 indicate that the effect of repetition neurons to prevent the repetition problem becomes smaller when using the larger language model. This result aligns with the experiment shown in the existing work Wang et al. (2024). In contrast, the activation of repetition neurons of Gemma-7B largely affects the repetitive outputs (Figure 12). These differences in performance show that there is room to be explored about the repetition neuron from broader viewpoints.

### v2

@@ 1 Introduction
While text generation with LLMs such as GPT-3 Brown et al. (2020) has been actively studied, the issue of repetition remains a fundamental challenge Li et al. (2023a); Ivgi et al. (2024). Specifically, repetition is particularly problematic under greedy generation, which is often used when reproducibility must be guaranteed Song et al. (2024).
Many researchers have tackled this problem by analyzing repetition Fu et al. (2021); Xu et al. (2022) and developing techniques to mitigate repetitive outputs Keskar et al. (2019); Shirai et al. (2021); Zhu et al. (2023); Li et al. (2023a). Some works specifically focus on attention heads, such as induction heads, framing repetition as a key mechanism for in-context learning Olsson et al. (2022); Bansal et al. (2023); Crosbie and Shutova (2024). However, the internal mechanisms of generative models that produce repetitive outputs remain insufficiently explored Vaidya et al. (2023); Wang et al. (2024).
We focus on the neurons of Transformer language models Vaswani et al. (2017); Geva et al. (2021); Dai et al. (2022); Chen et al. (2024) that detect repetition in inputs and trigger repetitive outputs in text generation. We refer to these neurons as “repetition neurons” following Wang et al. (2024). We hypothesize that the repetition neuron is a “skill neuron” Radford et al. (2017); Wang et al. (2022) that prompts the model to generate repetition as a task of copying the previous context, akin to “task vectors” Hendel et al. (2023) found in in-context learning Brown et al. (2020); Yan et al. (2024).
@@
We inspected repetition neurons in three English and one Japanese pre-trained language model. Our experimental results show that repetition neurons appear in both intermediate and final layers (§3.2). Furthermore, we demonstrate that deactivating these neurons suppresses the output probabilities of repeated tokens (§4.1), while activating them increases these probabilities (§4.2) [^1] Code for our experiments is available at https://github.com/tatHi/repetition_neuron. In addition, we highlight the relationship between repetition neurons and induction heads (§5).
@@ 3.2 Observation of Repetition Neuron
Figure 3 presents the distribution of repetition neurons across different layers. The last layer contains the largest number of repetition neurons in all models, while a secondary peak appears in the intermediate layers. This suggests the existence of two types of repetition neurons: those that detect repeating patterns in the intermediate layers and those that drive the model to replicate previous contexts in the uppermost layer. The presence of repetition neurons in both the final and intermediate layers aligns with previous findings that task-specific neurons tend to reside in higher layers Wang et al. (2022), and task-related parameters and hidden states are often found in intermediate layers Hendel et al. (2023); Merullo et al. (2024). Figure 9(a) and 9(b) in Appendix E show that the location patterns of the repetition neurons remain consistent across variations in hyperparameters $|X|$ and $r$.
@@ 4 Intervention to Repetition Neurons
[…]
@@ 4.1 Preventing Repetition
Result: Figure 4 shows the number of samples containing repetition after deactivating varying numbers of repetition neurons (solid lines) compared to randomly selected neurons (dashed lines). As the figure demonstrates, deactivating the repetition neurons effectively reduces the number of samples with repetition compared to deactivating randomly selected neurons. This result confirms that the repetition neurons identified by our method are indeed responsible for causing the repetition problem. We observed that deactivating repetition neurons reduces the number of samples with repetition by up to 25% (and by as much as 35% with optimal hyperparameter settings, as shown in Figures 10(a) and 10(b)). This suggests that roughly 30% of the repetition problem can be attributed to the repetition neurons. Table 3 in §C provides an example where repetition was successfully suppressed, illustrating that the generated text remains grammatically coherent despite neuron intervention. Besides, the perplexity is not largely damaged by deactivating the repetition neurons, as shown in Figure 8(a), which supports the coherency of the performance quantitatively. This confirms that the repetition neurons are specifically responsible for triggering repetition.
@@ 4.2 Invoking Repetition
5 Comparison with Induction Heads
Figure 6: Frequency of reputation neurons (lines), induction heads (colored bars), and self-finding heads (edged bars) for repetition over three English models.
Figure 7: Examples of attention heat-maps for induction and self-finding heads capturing repetition inputs: “a b c d e c d e c d e”. The expected next token is “c”.
Several works in in-context learning have examined how attention heads, particularly induction heads  Olsson et al. (2022); Bansal et al. (2023); Crosbie and Shutova (2024), exhibit repetitive behaviors. These studies explore how in-context learning performance relates to an LLM’s ability to copy patterns in synthetically generated repeating sequences. Building on these insights, we focus on a neuron-based analysis of repetition actually generated by the LLMs themselves. In this section, we compare the “repetition neurons,” “induction heads,” and “self-finding heads” derived from the same repeating texts using three LLMs: Gemma-2B, Pythia-2.8B, and LLaMA-3.2-3B.
Figure 6 shows the distribution of the repetition neurons (the same results as in Figure 3), induction heads, and self-finding heads over layers. In our analysis, we define “induction heads” as heads attending to repeating tokens that are to be generated after the current input token (Figure 7a). We also define “self-finding heads” as heads attending to the repeating token identical to the input token (Figure 7b). We identified a head as induction or self-finding if its total attention score for the target tokens that appear after the second repeating position exceeds 0.5. We then summarize their layer positions to see how these heads align with repetition neurons.
The observed behavior varies by model architecture. As shown in Figure 6a, Gemma-2B’s repetition neurons share two peaks with the induction heads, and one of these peaks is also shared by self-finding heads. This suggests that certain repetition neurons are activated in response to both induction and self-finding heads capturing repetition. However, the highest induction-head peak (layer 14 of 18) does not coincide with the highest repetition-neuron peak (layer 18 of 18).
Figures 6b and 6c present a different pattern for Pythia and LLaMA, where we do not observe a strong alignment between repetition neurons and induction heads. Nevertheless, similar to Gemma-2B, some peaks in the early layers of repetition neurons correspond to peaks of self-finding heads. This suggests that repetition neurons respond to self-finding patterns in earlier layers and take on different roles in later layers.
Overall, this comparison among repetition neurons, induction heads, and self-finding heads reveals coordinated interactions while showing their distinct roles in detecting and invoking repetition.
@@ 6 Conclusion
  • Beyond the neuron-based and head-based analysis (§5), can we find any other specific circuit in the LLMs’ calculation when outputting the repetitive texts?
@@ Appendix D Perplexity with Intervention
Figure 8(a) and 8(b) show the changes in perplexity based on different numbers of intervened repetition neurons. We used the test split of WikiText-2 for all models, including LLM-jp-1.8B, which was fine-tuned on Japanese corpora. As shown in Figure 8(a), the performance degradation caused by deactivating repetition neurons is relatively moderate. Although the impact on perplexity is smaller when intervening on randomly sampled neurons, considering that GPT-2’s perplexity on the same corpus was 18.34 Radford et al. (2019), the degradation caused by deactivating repetition neurons is acceptable. This suggests that repetition neurons do not significantly affect the generation of normal texts that do not contain repetition.
@@ Appendix E Ablation Study
Figure 9(a) and 9(b) show the location of the repetition neuron on the Transformer layers (§3.2). As shown in the figure, the size of the dataset $|X|$ does not have large effect on the distribution, which means we can obtain the similar set of repetition neurons both with smaller and larger sizes of datasets. On the other hand, $r$ has an effect on the distribution to some degree. For the case of Gemma-2B, we can obtain roughly a similar tendency with $15\leq r$.
@@
Figure 10(a) and 10(b) show the difference in the performance for the experiment about deactivating repetition neurons (§4.1). The experimental result indicates that we can obtain the more reduction effect with the larger number of deactivated repetition neurons. In contrast, the larger $r$ cause the decrease of the effect to reduce the repetition with larger number of deactivated neurons. Figure 10(b) suggests that $r=10$ or $r=15$ leads to the largest effect on controlling the repetitive generation.
@@
The experimental results for the effect of various hyperparameters on the setting with activating repetition neurons in Figure 11(a) and 11(b) show similar trends. The larger $|X|$ leads to the larger number of repetitive texts while $r=10$ or $r=15$ has the largest effect on the repetitive generation. These results could be an important clue to investigate the relation between the length of repetitive phrases and the repetition neurons in the future work.
@@ Appendix F Comparison between Two Model Sizes
Figure 13 and 14 show the experimental results of the two experiments with deactivating and activating the repetition neurons, respectively. The results of Figure 13 indicate that the effect of repetition neurons to prevent the repetition problem becomes smaller when using the larger language model. This result aligns with the experiment shown in the existing work Wang et al. (2024). In contrast, the activation of repetition neurons of Gemma-7B largely affects the repetitive outputs (Figure 14). These differences in performance show that there is room to be explored about the repetition neuron from broader viewpoints.